QUESTION # 1

Use the Student's t distribution to find tc for a 0.90 confidence level when the sample is 5. (Round your answer to three decimal places.)

QUESTION #2

Do you want to own your own candy store? Wow! With some interest in running your own business and a decent credit rating, you can probably get a bank loan on startup costs for franchises such as Candy Express, The Fudge Company, Karmel Corn, and Rocky Mountain Chocolate Factory. Startup costs (in thousands of dollars) for a random sample of candy stores are given below. Assume that the population of x values has an approximately normal distribution.

98 174 134 99 75 94 116 100 85

(a) Use a calculator with mean and sample standard deviation keys to find the sample mean startup cost x and sample standard deviation s. (Round your answers to one decimal place.)

x = __________ thousand dollars
s = ____________ thousand dollars


(b) Find a 90% confidence interval for the population average startup costs μ for candy store franchises. (Round your answers to one decimal place.)

lower limit     _________ thousand dollars
upper limit     __________thousand dollars
Answer #1

1)

sample size, n = 5
degrees of freedom, df = n - 1 = 4

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, tc = t(α/2, df) = 2.132


2)

a)

x =108.3
s = 29.9

b)

sample mean, xbar = 108.3
sample standard deviation, s = 29.9
sample size, n = 9
degrees of freedom, df = n - 1 = 8

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, tc = t(α/2, df) = 1.86


ME = tc * s/sqrt(n)
ME = 1.86 * 29.9/sqrt(9)
ME = 18.538

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (108.3 - 1.86 * 29.9/sqrt(9) , 108.3 + 1.86 * 29.9/sqrt(9))
CI = (89.8 , 126.8)


lower limit = 89.8
Upper limit = 126.8